OpenStudy (anonymous):

The population of a town was 6,000 in 1990 and grew to 9,000 in 2000. Assume that the population will continue to grow exponentially at the same rate. What will the population of the town be in 2010? Please help :O

OpenStudy (anonymous):

this is an exponential growth problems. just use the equation for exponential growth

OpenStudy (anonymous):

What is the equation?

OpenStudy (anonymous):

something like y = a(1 + b)x ?

OpenStudy (anonymous):

A(t))=Ainitial*e^k*t

OpenStudy (anonymous):

How do I figure out the rate of growth? Will you write a key corresponding the the letters in the equation as you typed it?

OpenStudy (anonymous):

A initial is 6000, A(t) 9000, e is constant 2.71..., k is a variable you find, t is the change it time( subtract final year minus initial)

OpenStudy (anonymous):

would you lke me to solve?

OpenStudy (anonymous):

So.. 9000=6000*2.71^k*20 and solve for k? Lol, sorry I'm so slow.. that isn't right is it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k=-2.5982 so now I just plug that back in for k?

OpenStudy (anonymous):

now read the question again

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

the question asks for the population in 2010, now you have to find final population A, A intial can be 6000 or 9000, time would be (2010 - intial time(depends on which initial value you use)) and you have k, solve A in A=Aintial*e^kt

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

a=9000*2.71^(-2.5982*20)

OpenStudy (anonymous):

there only errors now are math ones ^^.

OpenStudy (anonymous):

I'm getting weird answers form my calculator...

OpenStudy (anonymous):

value?

OpenStudy (anonymous):

my k is .000152

OpenStudy (anonymous):

sry its .1046

OpenStudy (anonymous):

i get 48665 what do u get?

OpenStudy (anonymous):

hah, 9027, fml

OpenStudy (anonymous):

maybe i should look at it again in the morning

OpenStudy (anonymous):

yeah rest ez man

OpenStudy (anonymous):

Thanks bro